OpenStudy (anonymous):

HELP PLEASE WILL GIVE FAN AND MEDAL AND I WILL LOVE YOU FOREVER (that sounded weird) BUT I WILL PLEASE I AM BEGGING YOU IT IS ALGEBRA PLEASE KNOW WHAT YOU ARE DOING

OpenStudy (anonymous):

ill try to help

OpenStudy (anonymous):

@dan815 @Data_LG2 @Conqueror @anybody lol

OpenStudy (anonymous):

ok @rubyredgirl20

OpenStudy (anonymous):

The expression below shows the population of butterflies in a forest after 4x years: f(x) = 40(1.04)4x Which of the following is an equivalent function? f(x) = 40(4.1600)x f(x) = 480(1.04)x f(x) = 40(1.1699)x f(x) = 480(4.1600)x

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

I dont think I can give an explanation though

OpenStudy (anonymous):

i think its b but im not completely sure.. :/

OpenStudy (anonymous):

im not there yet lol

OpenStudy (anonymous):

ok

OpenStudy (theraggedydoctor):

A. It's A. You don't have to love me forever, though. xD

OpenStudy (anonymous):

HAHAHA why is it a though ? @TheRaggedyDoctor

OpenStudy (theraggedydoctor):

My help is free :D I mean, if you insist, people loving me forever sounds pretty cool.

OpenStudy (anonymous):

are you sure, because i cannot afford to get any wrong

OpenStudy (theraggedydoctor):

Positive.

OpenStudy (anonymous):

ok thanks do you mind helping me w. a couple more ?

OpenStudy (theraggedydoctor):

I guess so, :P

OpenStudy (anonymous):

An expression is shown below: square root of 18 plus square root of 2 Which statement is true about the expression? It is rational and equal to 3. It is rational and equal to 4. It is irrational and equal to 3 multiplied by the square root of 2. It is irrational and equal to 4 multiplied by the square root of 2.

OpenStudy (theraggedydoctor):

D.

OpenStudy (anonymous):

are you sure i pretty sure it is b ?

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (theraggedydoctor):

Yep, I'm like 94.6% sure

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

The function f(x) = 800(1.09)3x shows the growth due to interest that Jim earns on an investment. The steps to calculate an equivalent function are shown below: f(x) = 800(1.09)3x [Step 1] f(x) = 800(1.093)x [Step 2] f(x) = 800(1.09 • 3)x [Step 3] f(x) = 800(3.27)x [Step 4] Which of the following is the first incorrect step? Step 1 Step 2 Step 3 Step 4

OpenStudy (theraggedydoctor):

Step 2.

OpenStudy (anonymous):

The number of bacteria in a sample increased by a factor of 2 every minute from 9:05 a.m. to 9:55 a.m. The function below shows the number of bacteria in the sample f(x) after x minutes from 9:05 a.m. : f(x) = 10(2)x Which of the following is a reasonable domain for the function? 0 ≤ x ≤ 10 5 ≤ x ≤ 55 0 ≤ x ≤ 50 All positive integers

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (theraggedydoctor):

0 ≤ x ≤ 50

OpenStudy (anonymous):

Sophia invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of her investment after x years: f(x) = 500(1.05)x What was the average rate of change of the value of Sophia's investment from the second year to the fourth year? 14.13 dollars per year 28.25 dollars per year 50.00 dollars per year 56.50 dollars per year

OpenStudy (deana99):

it is A

OpenStudy (theraggedydoctor):

I'd actually say B...

OpenStudy (deana99):

why

OpenStudy (anonymous):

which one is it ?

OpenStudy (theraggedydoctor):

I'm pretty darn sure it's B. I checked my work. @deana99 how did you get A?

OpenStudy (anonymous):

So what is it @TheRaggedyDoctor and @deana99 please cofrim the answer .

OpenStudy (deana99):

its B i did something wrong

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

A music competition on television had five elimination rounds. After each elimination, only half of the contestants were sent to the next round. The table below shows the number of contestants in each round of the competition: x 1 2 3 4 5 f(x) 64 32 16 8 4 Compute the average rate of change of f(x) from x = 3 to x = 5 and describe what it represents. -6 contestants per round, and it represents the number of contestants who will reach the final round -28 contestants per round, and it represents the number of contestants who will reach the final round -6 contestants per round, and it represents the average rate at which the number of contestants decreased from the third round to the fifth round -28 contestants per round, and it represents the average rate at which the number of contestants decreased from the third round to the fifth round

OpenStudy (anonymous):

this one is either a or c

OpenStudy (theraggedydoctor):

C

OpenStudy (deana99):

its A

OpenStudy (anonymous):

will you please close this question and PM each other? im tired of getting my notifications blown up.. srry if that sounded rude, but im waiting for a certain notification and this is kind getting annoying

OpenStudy (anonymous):

I will post another tab and tag you guy s

OpenStudy (deana99):

ok

OpenStudy (anonymous):

alright, thanks :)